OpenStudy (anonymous):

The midpoint of the segment joining points (a, b) and ( j, k) is:

OpenStudy (anonymous):

here are some options

OpenStudy (jim766):

to find the mid point add the x's and div by 2 add the y's and div by 2 (a, b) (j, K) ----- a+j , k+b --- --- 2 2

OpenStudy (jim766):

make sense?

OpenStudy (anonymous):

ya thanks

OpenStudy (jim766):

yw

OpenStudy (anonymous):

can you help with this one?

OpenStudy (anonymous):

Point T is the midpoint of JH . The coordinate of T is (0, 4) and the coordinate of J is (0, 2). The coordinate of H is:

OpenStudy (anonymous):

a.(0, 6) b.(0, 3) c.(0, 7) d.(0, 11)

OpenStudy (jim766):

in this one the give you the mid point....so we need to do the opposite of before. start with the mid point and instead of div by 2, mult by 2 (0, 4) * 2 = (0, 8 ) now subtract the other end - (0, 2 ) ------- =

OpenStudy (anonymous):

so it A right?

OpenStudy (jim766):

perfect!

OpenStudy (anonymous):

what about this one Point (-4, 3) lies in Quadrant a.I b.II c.III d.IV

OpenStudy (jim766):



OpenStudy (jim766):

do you know how to graph the point?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so its B

OpenStudy (jim766):

good!

OpenStudy (anonymous):

Point (6, 0) lies on the a.x-axis b.y-axis c.line y = x

OpenStudy (anonymous):

it s x-axis right

OpenStudy (anonymous):

???

OpenStudy (jim766):

right! x y (6, 0) over right 6, and zero up or down

OpenStudy (anonymous):

what about this one? Any line with no slope is parallel to the a.x-axis b.y-axis c.line y = x

OpenStudy (jim766):

if it has no slope, that means it is flat... and horizontal line which axis is horizontal?

OpenStudy (anonymous):

X

OpenStudy (jim766):

right!

OpenStudy (anonymous):

If the point (a,3) lies on the graph of the equation 5x + y = 8, then a= a.1 b.-1 c.-7

OpenStudy (jim766):

they have given you a y x y (a, 3) to find the other just sub in the 3 for y 5x + 3 = 8, now solve for x

OpenStudy (anonymous):

x=1 right

OpenStudy (jim766):

great!

OpenStudy (anonymous):

If the equation of a circle is (x + 5)^2 + (y - 7)^2 = 36, its center point is a.(5, 7) b.(-5, 7) c.(5, -7)

OpenStudy (jim766):

in this take the opposite of each in the ( )'s r = (-5, 7)

OpenStudy (anonymous):

If the equation of a circle is (x + 5)^2 + (y - 7)^2 = 36, its radius is a.6 b.16 c.36

OpenStudy (jim766):

take the sq root of what the equation = ____

OpenStudy (anonymous):

i dont get it???

OpenStudy (jim766):

(x + 5)^2 + (y - 7)^2 = 36 it = 36, find the sq root of 36

OpenStudy (anonymous):

sq root of 36 is 6 right

OpenStudy (anonymous):

???

OpenStudy (jim766):

yes that is the radis

OpenStudy (jim766):

6

OpenStudy (anonymous):

so 6 is the answer right?

OpenStudy (jim766):

yes,

OpenStudy (anonymous):

If the equation of a circle is (x - 2)^2 + (y - 6)^2 = 4, the center is point (2, 6). True False

OpenStudy (jim766):

what do you think?

OpenStudy (anonymous):

true

OpenStudy (jim766):

yep

OpenStudy (jim766):

it was nice helping you , but gotta run

OpenStudy (anonymous):

okay thanks alot